Reid Company is considering the production of a new product. The expected variable cost is $30 per unit. Annual fixed costs a6
Answer #1

Answer

  • Requirements asked

a

Break even in units

                             17,500

=875000/(80-30)

Break even point in dollars

$1,400,000

=17500*80

b

Contribution margin per unit

$50

=80-30

Break even in units

                              17,500

=875000/50

Break even point in dollars

$1,400,000

=17500*80

c

CM Ratio

62.5%

=50/80

Break even in units

                              17,500

=+(875000/62.5%)/80

Break even point in dollars

$1,400,000

=875000/62.5%